OpenStudy (anonymous):

Determine the probability that the spinner lands on red. 4 red 2green 2 yellow 0.5 0.25 0.33 1.5 0.67

OpenStudy (anonymous):

0.5

OpenStudy (anonymous):

4/8=1/2=0.5

OpenStudy (anonymous):

thanks